erify that the following cquations are integrable and determine their solutions.
x(y^2-a^2 )dx+y(x^2-t^2 )dy-z(y^2-α^2 )dz=0, where α is constant
1
Expert's answer
2021-05-07T11:36:42-0400

The given equation is-

"x(y^2-a^2 )dx+y(x^2-t^2 )dy-z(y^2-\u03b1^2 )dz=0,"


Given equation is of the form-

"Pp+Qq+Rr=0"


The "\dfrac{dx}{P}=\dfrac{dy}{Q}=\dfrac{dz}{R}"


So,"\dfrac{dx}{x(y^2-a^2)}=\dfrac{dy}{y(x^2-t^2)}=\dfrac{dz}{z(y^2-\alpha^2)}"



Taking first two terms-

 "\dfrac{dx}{x(y^2-a^2)}=\dfrac{dy}{y(x^2-t^2)}"



"\Rightarrow \dfrac{(x^2-t^2)dx}{x}=\dfrac{(y^2-a^2)dy}{y}"


Integrating both the sides-


"\dfrac{x^2}{2}-t^2lnx=\dfrac{y^2}{2}-a^2lny+c_1"


"\Rightarrow \dfrac{x^2}{2}-t^2lnx-\dfrac{y^2}{2}+a^2lny=c_1~~~~~~~~~-(1)"


Taing last two terms-


"\dfrac{dy}{y(x^2-t^2)}=\dfrac{dz}{z(y^2-\alpha^2)}"



"\Rightarrow \dfrac{(y^2-\alpha^2)dy}{y}=\dfrac{(z^2-t^2)dz}{z}"


Integrating Both the sides-


"\dfrac{y^2}{2}-\alpha^2lny=\dfrac{z^2}{2}-t^2lnz+c_2"


"\Rightarrow \dfrac{y^2}{2}-\alpha^2lny-\dfrac{z^2}{2}+t^2lnz=c_2"



Hence The solution is-


"\phi(c_1,c_2)=0"


"\phi (\dfrac{x^2}{2}-t^2lnx-\dfrac{y^2}{2}+a^2lny, \dfrac{y^2}{2}-\alpha^2lny-\dfrac{z^2}{2}+t^2lnz)=0"